# An Objective Improvement Approach

### v2

@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f)$
These steps can be identified in Algorithm 2. The objective function is fixed at the beginning (Line 1). It simply maximises the sum of the values of all the vertices, and thus implicitly the value of all vertices. (1) The strategy of one of the two player (for convention the maximiser – if we want to instead fix the strategy of the minimiser, we would have to minimise the sum of the values for all vertices instead) is initialised at Line 2. (2) Before the computation of the valuation at Line 5, the set of constraints is updated accordingly to the selected strategy of the maximiser (Line 4). At this point, if the strategy is optimal w.r.t. the current valuation, then the algorithm ends. Otherwise, (3) there exist an improvement, which is applied at Line 7.
@@
Thus, in our novel symmetric approach the constraints remain while the objective is updated, in strategy improvement, instead, the objective remains while the constraints are updated. Moreover, the players and their strategies are treated quite differently in strategy improvement algorithms: while the candidate strategy of the maximiser results in making the inequations of the selected edges sharp (and dropping all other inequations of maximiser edges), the optimal counter strategy is found by minimising the objective. This is again in contrast to our novel symmetric approach, which treats both players equally.
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f_{\sigma})$
The inequations from $H$ are computed by a function called $\mathsf{Inequations}$ (Line 1) that, given the discounted game $\mathcal{G}$, returns the set made of one inequation per edge $e=(v,v^{\prime})\in E$, defined as follows:
@@
We put a further restriction on $\mathsf{val}$, returned by the call $\mathsf{LinearProgramming}$, in that we require it to be the solution to a basis $\mathbf{b}$ in $H$. Such a basis consists of $|V|$ inequations that are satisfied sharply (again, as equations), such that, accordingly to the definition of basis, these $|V|$ equations uniquely define the values of all vertices. We denote with $\mathsf{val}_{\mathbf{b}}$ the evaluation of $\mathbf{b}$. As a consequence, the call $\mathsf{LinearProgramming}(H,f_{\sigma})$ to some linear programming algorithm returns a valuation $\mathsf{val}$ of the vertices that minimise $f_{\sigma}$ while satisfying $H$ (Line 5), and for convenience require this valuation to also be $\mathsf{val}_{\mathbf{b}}$ for some basis $\mathbf{b}$ of $H$. (Note that the simplex algorithm, for example, only uses valuations of this form in every step.) We call this valuation a valuation associated to $\sigma$.
@@
If, for a solution $\mathsf{val}$ of $H$, $f_{\sigma}(\mathsf{val})=0$ holds, then, for every vertex $v\in V$, the inequation $I_{(v,\sigma(v))}$ for the edge $(v,\sigma(v))$ is sharp, and $\mathsf{val}$ therefore defines strategies for both players – those defined by $\sigma$, for example.
Instead of checking whether $\mathsf{val}$ defines strategies of both players, we can use, alternatively, $f_{\sigma}(\mathsf{val})=0$ as a termination condition, as shown at Line 6 in Algorithm 3, since in this case $\sigma$ must define co-optimal strategies.
@@
The valuation $\mathsf{val}=\mathsf{val}_{\mathcal{G}}$ of the game is the unique solution of $H$ for which, for all vertices $v$, the inequation to (at least) one of the outgoing edges of $v$ is sharp, and the edges for which they are sharp describe co-optimal strategies. This valuation of the game is thus the only valuation that defines strategies for both players, which shows the second claim.
@@
The theorem above ensures that, in case the condition at Line 6 holds, the algorithm terminates and provides the value of the game that then allows us to infer optimal strategies of both players. Otherwise we have to improve the objective function and make another iteration of the while loop. At Line 7, $\mathsf{ChooseBetterStrategies}$ can be any procedure that, for $f_{\sigma}(\mathsf{val})\neq 0$, provides a pair of strategies $\sigma^{\prime}$ better than $\sigma$ as defined as follows.
@@
Better Strategy
A strategy $\sigma^{\prime}$ for both players is better than a strategy $\sigma$ if, and only if,
@@
Given a game $\mathcal{G}$, a strategy $\sigma^{\prime}$ and the corresponding objective function $f_{\sigma^{\prime}}$ computed by $\mathsf{ObjectiveFunction}(\mathcal{G},\sigma)$, the call $\mathsf{LinearProgramming}(H,f_{\sigma^{\prime}})$ computes the minimal value of $f_{\sigma^{\prime}}$. If this value is strictly lower than the minimal value for $f_{\sigma}$, computed by $\mathsf{LinearProgramming}(H,f_{\sigma})$, then $\sigma^{\prime}$ is better than $\sigma$.
@@ 5. Choosing Better Strategies
    If the current valuation $\mathsf{val}$ does not define a pair of strategies $\sigma$ for both players and has no local improvements, then a better strategy $\sigma^{\prime}$ can be found by applying only switches from and to edges that already have offset 00 (Section 5.2).
@@
    The improvement mentioned in the previous point can be found for special games – the sharp and improving games defined in Section 5.3 – by trying a single edge switch.
@@
As a small side observation, when using a simplex based technique to implement $\mathsf{LinearProgramming}$ at Line 5 of Algorithm 3, then the pivoting of the objective function from point (1) and the pivoting of the basis can be mixed (this will be discussed in Section 5.5).
@@ 5.1. Local Improvements
The simplest – and most common – case of creating better strategies $\sigma^{\prime}$ from a valuation for the objective $f_{\sigma}$ for a strategy $\sigma$ is to consider local improvements. Much like local improvements in strategy iteration approaches, local improvements consider, for each vertex $v$, a successor $v^{\prime}\neq\sigma(v)$, such that $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))<\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for the current valuation $\mathsf{val}$, which is optimal for the objective function $f_{\sigma}$.
To be more general, our approach does not necessarily require to select only local improvements, but it can work with global improvements, though we cannot see any practical use of choosing differently. For instance, if we treat the function as a global improvement approach, we can update the value for a vertex $v$ such that it increases by 1 and update the value of another vertex $v^{\prime}$ such that it decreases by 2. The overall value of the function will decrease, even if locally some components increased their value. Interestingly, this cannot be done with a strategy improvement approach, as it requires to always locally improve the value of each vertex when updating.
@@
The valuation $\mathsf{val}$ is, being an optimal solution for the objective $f_{\sigma}$, a solution to the system of inequations $H$. For a solution $\mathsf{val}^{\prime}$ which is optimal for $f_{\sigma^{\prime}}$, we thus have $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})\leq f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, which implies that $\sigma^{\prime}$ is better than $\sigma$ accordingly to Definition 4.1 of better strategy.
@@ 5.2. No Local Improvements
Note that $S_{\mathsf{val}}^{\sigma}$ is such a set, hence, an adequate set is easy to identify. We might, however, be interested in keeping the set small, and choosing the edges defined by $E_{\mathsf{val}}$ plus one outgoing edge for every vertex $v$ that does not have an outgoing edge in $E_{\mathsf{val}}$. The most natural solution is to choose the edge $(v,\sigma(v))\in E_{\mathsf{val}}^{\sigma}$ defined by $\sigma$ for each such vertex $v$.
@@
By Observation 5.2, $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E_{\mathsf{val}}^{\sigma},w,\lambda)$ is a DPG. Let $\mathsf{val}^{\prime}$ be the valuation of $\mathcal{G}^{\prime}$, and $\sigma^{\prime}$ be the strategies for the two players defined by it.
If $\mathsf{val}^{\prime}$ is also a valuation of $\mathcal{G}$, then we are done. However, this need not be the case, as the system of inequations $H^{\prime}$ for $\mathcal{G}^{\prime}$ is smaller than the set of inequations $H$ for $\mathcal{G}$, so $\mathsf{val}^{\prime}$ might violate some of the inequations that are in $H$, but not in $H^{\prime}$. Given that $\mathsf{val}^{\prime}$ is a valuation for $\mathcal{G}^{\prime}$, it satisfies all inequations in $H^{\prime}$. Moreover, since $\mathsf{val}$ also satisfies all inequations of $H^{\prime}$, it follows that the same inequations hold for every convex combination of $\mathsf{val}$ and $\mathsf{val}^{\prime}$.
@@
We now have $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})>0$ and $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})=0$. For an optimal solution $\mathsf{val}^{\prime\prime}$ of $H$ for the objective $f_{\sigma^{\prime}}$, this provides $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})\leq f_{\sigma^{\prime}}(\mathsf{val}_{\varepsilon})<f_{\sigma}(\mathsf{val})$.
Therefore $\sigma^{\prime}$ is better than $\sigma$.
When using the most natural choice, $E_{\mathsf{val}}^{\sigma}=E_{\mathsf{val}}\cup\{(v,\sigma(v))\mid v\in V\}$, this translates in keeping all transitions, for which the offset is not 00, while changing some of those, for which the offset already is 00. This is a slightly surprising choice, since to progress one has to improve on the transitions whose offset is positive, and ignore those with offset 0.
@@ 5.3. Games with Efficient Objective Improvement
We now describe two conditions that, when simultaneously met, allow us to efficiently find better strategies: that games are sharp and improving.
@@
Sharp Game
A game $\mathcal{G}$ is called sharp for the valuation $\mathsf{val}$ if, and only if, $\mathsf{val}$ satisfies exactly $|V|$ inequations $H$ computed by $\mathsf{Inequations}(\mathcal{G})$.
@@
Improving Game
A game $\mathcal{G}$ is called $\sigma$-improving for an objective function $f_{\sigma}$ defined by a joined strategy $\sigma$ and a non-optimal valuation $\mathsf{val}$ defined by a basis for $f_{\sigma}$ if, and only if, there is a single basis change that leads to a better strategy.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is (at least) one neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
Such a strategy $\sigma^{\prime}$ is better than $\sigma$, and it can be selected in a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
@@
We apply Lemma 4 for $E_{\mathsf{val}}^{\sigma}$ that, we recall, corresponds to $E_{\mathsf{val}}\cup\{(v,\sigma(v)\mid v\in V\}$, and use the resulting better strategy $\sigma^{\prime}$ from this set $E^{\sigma}_{\mathsf{val}}$. Note that this means $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$, as there are no local improvements.
We now consider an arbitrary sequence of evaluations $\mathsf{val}_{0},\mathsf{val}_{1},\ldots,\mathsf{val}_{n}$ obtained by calls to the simplex algorithm on the linear program defined by the constraints $H$ obtained from $\mathcal{G}$. The evaluations correspond to adjacent corners of the polytope on which the simplex algorithm moves. The sequence starts from $\mathsf{val}_{0}=\mathsf{val}$ and ends with $\mathsf{val}_{n}=\mathsf{val}^{\prime}$. The value of the new objective function $f_{\sigma^{\prime}}$ can only decrease along this path between the corners of the polytope and $\mathsf{val}^{\prime}$ minimises $f_{\sigma^{\prime}}$.
Note that, since $\sigma^{\prime}$ is better than $\sigma$ by Lemma 4, we have that $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$. Moreover, the sharpness of $\mathcal{G}$ implies that $\mathsf{val}_{1}\neq\mathsf{val}_{0}$ and, since $\mathcal{G}$ is improving (and therefore $\sigma^{\prime}$-improving), this provides that $f_{\sigma^{\prime}}(\mathsf{val}_{1})<f_{\sigma^{\prime}}(\mathsf{val}_{0})$. Finally, we have already established $f_{\sigma^{\prime}}(\mathsf{val}_{0})=f_{\sigma}(\mathsf{val}_{0})$.
Thus, when only applying a single basis change, we move to a fresh value $\mathsf{val}^{\prime\prime}=\mathsf{val}_{1}$ such that $f_{\sigma^{\prime}}(\mathsf{val}_{1})<f_{\sigma}(\mathsf{val})$ for some $\sigma^{\prime}$.
Note that $\sigma^{\prime}$ was supplied by Lemma 4, so that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds.
While we can restrict the selection of $\sigma^{\prime}$ to the strategies that comply with the restriction $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$, there is no particular reason for doing so; as soon as we have a neighbouring valuation $\mathsf{val}^{\prime}$, we can identify a pair of strategies $\sigma^{\prime}$ for which $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})$ is minimal, and select $\sigma^{\prime}$ if $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ holds.

### v3

@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f)$
These steps can be identified in Algorithm 2. The objective function is fixed at the beginning (Line 1) and corresponds to the sum of the values of all the vertices. The algorithm simply maximises it [^5] The optimal minimiser strategy chooses edge that minimises the value for every minimiser vertex. In the linear program, this is revlected by an inequeation for every outgoing edge, like the inequation $\mathsf{val}(a)\leq 1+\mathsf{val}(a)$ from the running example, and then maximises the value, which forces the selected value to be the minimal value such that one such inequation is sharp. by searching for a solution that maximises the value of all vertices. (1) The strategy of one of the two players (for convention the maximiser – if we want to instead fix the strategy of the minimiser, we would have to minimise the sum of the values for all vertices instead) is initialised at Line 2. (2) Before the computation of the valuation at Line 5, the set of constraints is updated accordingly to the selected strategy of the maximiser (Line 4). At this point, if the strategy is optimal w.r.t. the current valuation, then the algorithm ends. Otherwise, (3) there exists an improvement, which is applied at Line 7.
@@
Thus, in our novel symmetric approach, the constraints remain while the objective is updated; in strategy improvement, instead, the objective remains while the constraints are updated. Moreover, the players and their strategies are treated quite differently in strategy improvement algorithms: while the candidate strategy of the maximiser results in making the inequations of the selected edges sharp (and dropping all other inequations of maximiser edges), the optimal counter strategy is found by minimising the objective. This is again in contrast to our novel symmetric approach, which treats both players equally.
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f_{\sigma})$
The inequations from $H$ are computed by a function called $\mathsf{Inequations}$ (Line 1) that, given the discounted game $\mathcal{G}$, returns the set made up of one inequation per edge $e=(v,v^{\prime})\in E$, defined as follows:
@@
We put a further restriction on $\mathsf{val}$, returned by the call $\mathsf{LinearProgramming}$, in that we require it to be the solution to a basis $\mathbf{b}$ in $H$. Such a basis consists of $|V|$ inequations that are satisfied sharply (again, as equations), such that, according to the definition of basis, these $|V|$ equations uniquely define the values of all vertices. We refer to this unique set of values as the evaluation of $\mathbf{b}$, denote by $\mathsf{val}_{\mathbf{b}}$.
The call $\mathsf{LinearProgramming}(H,f_{\sigma})$ to some linear programming algorithm returns a valuation $\mathsf{val}$ of the vertices that minimises $f_{\sigma}$ while satisfying $H$ (Line 5); for convenience, we require this valuation to also be the evaluation $\mathsf{val}_{\mathbf{b}}$ for some basis $\mathbf{b}$ of $H$. (Note that the simplex algorithm, for example, only uses valuations of this form in every step.) We call this valuation a valuation associated to $\sigma$.
@@
If, for a solution $\mathsf{val}$ of $H$, $f_{\sigma}(\mathsf{val})=0$ holds, then, for every vertex $v\in V$, the inequation $I_{(v,\sigma(v))}$ for the edge $(v,\sigma(v))$ is sharp and $\mathsf{val}$ therefore defines strategies for both players – those defined by $\sigma$, for example.
Instead of checking whether $\mathsf{val}$ defines the strategies of both players, we can use, alternatively, $f_{\sigma}(\mathsf{val})=0$ as a termination condition, as shown at Line 6 in Algorithm 3, since in this case $\sigma$ must define co-optimal strategies.
@@
The valuation $\mathsf{val}=\mathsf{val}_{\mathcal{G}}$ of the game is the unique solution of $H$ for which, for all vertices $v$, the inequation to (at least) one of the outgoing edges of $v$ is sharp. This case implies that the edges for which they are sharp describe co-optimal strategies. This solution of $H$ is thus the only one that defines strategies for both players, which shows the second claim.
@@
The theorem above ensures that, in case the condition at Line 6 holds, the algorithm terminates and provides the value of the game that then allows us to infer optimal strategies of both players. Otherwise, we have to improve the objective function and make another iteration of the while loop. At Line 7, $\mathsf{ChooseBetterStrategies}$ can be any procedure that, for $f_{\sigma}(\mathsf{val})\neq 0$, provides a joint strategy $\sigma^{\prime}$ better than $\sigma$ as defined below.
@@
(Better Strategy) A joint strategy $\sigma^{\prime}$ for both players is better than a joint strategy $\sigma$ if, and only if,
@@
Given a game $\mathcal{G}$, a strategy $\sigma^{\prime}$ and the corresponding objective function $f_{\sigma^{\prime}}$ computed by $\mathsf{ObjectiveFunction}(\mathcal{G},\sigma^{\prime})$, the call $\mathsf{LinearProgramming}(H,f_{\sigma^{\prime}})$ computes the minimal value of $f_{\sigma^{\prime}}$. If this value is strictly lower than the minimal value for $f_{\sigma}$, computed by $\mathsf{LinearProgramming}(H,f_{\sigma})$, then $\sigma^{\prime}$ is better than $\sigma$.
@@ 5. Choosing Better Strategies
    If the current valuation $\mathsf{val}$ does not define a pair of strategies $\sigma$ for both players and has no local improvements, then a better strategy $\sigma^{\prime}$ can be found applying only switches from and to edges that already have offset 00 (Section 5.2).
@@
    The improvement mentioned in the previous point can be found for special games – the sharp and improving games defined in Section 5.3 – by trying a single-edge switch.
@@
As a small side observation, when using a simplex-based technique to implement $\mathsf{LinearProgramming}$ at Line 5 of Algorithm 3, then the pivoting of the objective function from point (1) and the pivoting of the basis can be mixed (this will be discussed in Section 5.5).
@@ 5.1. Local Improvements
The simplest and most common case of creating better strategies $\sigma^{\prime}$ from a valuation for the objective $f_{\sigma}$ for a strategy $\sigma$ is to consider local improvements. Much like local improvements in strategy iteration approaches, local improvements consider, for each vertex $v$, a successor $v^{\prime}\neq\sigma(v)$, such that $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))<\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for the current valuation $\mathsf{val}$, which is optimal for the objective function $f_{\sigma}$.
To be more general, our approach does not necessarily require one to select only local improvements, but it can work with global improvements, though we cannot see any practical use of choosing differently. For example, if we treat the function as a global improvement approach, we can update the value of a vertex $v$ so that it increases by 1 and update the value of another vertex $v^{\prime}$ so that it decreases by 2. The overall value of the function will decrease even if locally some components increased their value. Interestingly, this cannot be done with a strategy improvement approach, as it requires one to always locally improve the value of each vertex when updating.
@@
The valuation $\mathsf{val}$ is, being an optimal solution for the objective $f_{\sigma}$, a solution to the system of inequations $H$. For a solution $\mathsf{val}^{\prime}$ that is optimal for $f_{\sigma^{\prime}}$, we thus have $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})\leq f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, which implies that $\sigma^{\prime}$ is better than $\sigma$ according to Definition 4.1 of a better strategy.
@@ 5.2. No Local Improvements
Note that $S_{\mathsf{val}}^{\sigma}$ is such a set, and, therefore, an adequate set is easy to identify. However, we might be interested in keeping the set small and choosing the edges defined by $E_{\mathsf{val}}$ plus one outgoing edge for every vertex $v$ that does not have an outgoing edge in $E_{\mathsf{val}}$. The most natural solution is to choose the edge $(v,\sigma(v))\in E_{\mathsf{val}}^{\sigma}$ defined by $\sigma$ for each such vertex $v$.
@@
By Observation 5.2, $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E_{\mathsf{val}}^{\sigma},w,\lambda)$ is a DPG. Let $\mathsf{val}^{\prime}$ be the value of $\mathcal{G}^{\prime}$, and $\sigma^{\prime}$ be the strategies for the two players defined by it.
If $\mathsf{val}^{\prime}$ is also a solution of $\mathcal{G}$, then we are done. However, this need not be the case, as the set of inequations $H^{\prime}$ for $\mathcal{G}^{\prime}$ is smaller than the set of inequations $H$ for $\mathcal{G}$, so $\mathsf{val}^{\prime}$ might violate some of the inequations that are in $H$, but not in $H^{\prime}$. Given that $\mathsf{val}^{\prime}$ is a solution for $\mathcal{G}^{\prime}$, it satisfies all inequations in $H^{\prime}$. Moreover, since $\mathsf{val}$ also satisfies all inequations of $H^{\prime}$, it follows that the same inequations hold for every convex combination of $\mathsf{val}$ and $\mathsf{val}^{\prime}$.
@@
As there is no local improvement, we have $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))\geq\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for every vertex $v$. At the same time, as $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$, we have $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))=\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ or $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))=0$; together, this provides $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))=\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$. As this holds for every vertex, this provides $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$. Thus, we now have $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})>0$ and $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})=0$. For an optimal solution $\mathsf{val}^{\prime\prime}$ of $H$ for the objective $f_{\sigma^{\prime}}$, this provides $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})\leq f_{\sigma^{\prime}}(\mathsf{val}_{\varepsilon})<f_{\sigma}(\mathsf{val})$.
Therefore, $\sigma^{\prime}$ is better than $\sigma$.
When using the most natural choice, $E_{\mathsf{val}}^{\sigma}=E_{\mathsf{val}}\cup\{(v,\sigma(v))\mid v\in V\}$, this translates into keeping all transitions, for which the offset is not 00, while changing some of those for which the offset already is 00. This is a slightly surprising choice, since to progress one has to improve on the transitions whose offset is positive, and ignore those with offset 0.
@@ 5.3. Games with Efficient Objective Improvement
We now describe two conditions that, when simultaneously met, allow us to efficiently find better strategies: that the games are sharp and improving.
@@
(Sharp Game) A game $\mathcal{G}$ is called sharp for a valuation $\mathsf{val}$ if, and only if, $\mathsf{val}$ satisfies exactly $|V|$ inequations $H$ computed by $\mathsf{Inequations}(\mathcal{G})$ sharply. It is called sharp if no solution satisfies strictly more than $|V|$ inequations sharply.
@@
{obs}
Bases define valuations, and valuations define (sets of strategies): as we have used in Lemma 3, it is easy to identify for a valuation $\mathsf{val}$ the pair of strategies $\sigma$ (or the pairs of strategies if there are many) for a whom $f_{\sigma}(\mathsf{val})$ is smallest among all pairs of strategies; they are simply those pairs that minimise the offset for every vertex.
We therefore say that a value defines these pairs of strategies, and that a basis defines the pairs of strategies defined by the valuation it defines.
@@
(Improving Game) A game $\mathcal{G}$ is called $\sigma$-improving for an objective function $f_{\sigma}$ defined by a joined strategy $\sigma$ and a non-optimal valuation $\mathsf{val}$ defined by a basis for $f_{\sigma}$ if, and only if, there is a single basis change that leads to a better strategy (having a smaller value of $f_{\sigma}$) defined by the new basis.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG, $\sigma$ a strategy for both players that is not co-optimal, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is (at least) one neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
This strategy $\sigma^{\prime}$ can be selected in such a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
@@
By Lemma 4, we can obtain a better strategy $\sigma^{\prime}$ from transitions in $E_{\mathsf{val}}^{\sigma}$; for $\sigma^{\prime}$ therefore $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$ holds, and $\mathsf{val}$ is not optimal for $f_{\sigma}$.
Let $\mathbf{b}$ be a basis that defines $\mathsf{val}$. As $\mathcal{G}$ is improving, it is $\sigma^{\prime}$-improving, so that there is basis $\mathbf{b}^{\prime}$ that neighbours $\mathbf{b}$ and defines a valuation $\mathsf{val}^{\prime}$ with $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$.
We observe that any set $E^{\sigma}_{\mathsf{val}}$ can be selected, including the set $E_{\mathsf{val}}\cup\{(v,\sigma(v)\mid v\in V\}$, to select a better strategy $\sigma^{\prime}$ from.
While we can restrict the selection of $\sigma^{\prime}$ to the strategies that comply with the restriction $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$, there is no particular reason for doing so; as soon as we have a neighbouring valuation $\mathsf{val}^{\prime}$, we can identify a pair of strategies $\sigma^{\prime}$ for which $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})$ is minimal and select $\sigma^{\prime}$ if $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ holds.
Remark.
As an alternative to considering all neighbouring bases of $\mathbf{b}$, determining their valuation, and determining the objective function for $\sigma^{\prime}$, we can also determine the reduced cost for exchanging a basis based on the best improvement obtained by any function derived from $E^{\sigma}_{\mathsf{val}}$ and use this to locate an improvement.